Subject: Minutes of APO User's Committee Meeting, Feb 10, 1997

From: strauss@astro.Princeton.EDU

Submitted: Wed, 19 Feb 1997 11:52:05 -0500 (EST)

Message number: 126 (previous: 125, next: 127 up: Index)

   Minutes of APO 3.5m User's Committee Phone Conference
		  Monday, February 10, 1997

   Attending: Bruce Gillespie, Ed Turner (Chair), Michael Strauss
(taking minutes), Rene Walterbos, Alan Uomoto, Lew Hobbs, Jeff Secker.

Jeff Secker is taking over from Julie Lutz as Washington State
representative on the User's Committee.

  Agenda Items:
CCD cameras for the Guider and for the Echelle
Observations of Hale-Bopp with the 3.5m?
SPICam status, future, and user documentation
Targets of Opportunity programs
New DIS gratings
New chips for DIS?
Recent Pointing Problems
Recent Computer Glitches
Continued monitoring of the throughput
Instrument issues

  The 512x512 camera currently in use on the guider is needed by the
Echelle people for development of their guider.  We will get them a
different camera to work with, and so the conflict is resolved.  Indeed,
the new camera used at Nasmyth, from Photometrics, should be much
better.  We thus will be able to get rid of the current 512x512 camera
from Spectrasource, which is quite poor, by March-ish.  With the
improved camera, and the improved throughput of the telescope, this
new camera should go deep enough to find guide stars at any position
in the sky, at any position of the sky.  This is wonderful news!
The new camera being bought for the Echelle will be compatible with the
Nasmyth guider camera, and can be used as a spare.

Hale-Bopp: should we be doing anything?
  Comet Hale-Bopp is currently in the skies; it is perhaps a bit
embarrassing that APO does not have any specific plans to observe it.
  Uomoto: Paul Feldman will be launching a rocket from White Sands to
observe in mid-March; it would be good to get spectra at that time
(although it is only 15 degrees above the horizon at twilight at that
time).  Uomoto and Turner will discuss various observing
possibilities, including taking images with SPICAM.  If people have
further ideas about comet observing programs, they should talk to Turner.

SPICAM and its status:
  Documentation is starting to be put together on SPICAM performance,
specifications, etc.  Walterbos says that one of the NMSU people would
be happy to write up a manual, and do a proper throughput study.
  A filter wheel exists; they are now working on making it
controllable remotely.
  Can users yet ask for SPICAM?  There is not yet a hard timetable for
when it will be generally available.  The plan is that this camera
will replace DSC once it is fully commissioned, and it has very
similar characteristics (i.e., type of chip, field of view, pixel
size, capabilities, etc) as does DSC (more details wait on the
detailed documentation to be put together for the camera).  Thus
observers who wish to use SPICAM for second quarter should indicate on
their form whether DSC would be acceptable if SPICAM is not available
(and vice-versa).
  It is not clear whether the chip in SPICAM is thinned.  If it is
not, then (unlike DSC) SPICAM has very little response in the UV.

  Tim McKay says that the people at Yerkes might have interest in using DSC
there after it is decommissioned at APO.  He also points out that the
SDSS filter set currently in the DSC are not owned by APO, and will
leave with the instrument; we will need to acquire a new set of 3x3
SDSS filters for APO (about $5K).
  In any case, the filters that are available are documented on the
APO Web site.  SPICAM takes 3x3 filters; it is not known if it takes
2x2 filters as well.

  Jim Fowler and Bob Loewenstein are starting to think about developing
software to allow
SPICAM to be controlled through REMARK.  It will still be another year
before we start thinking about rewriting the basic observing software
from scratch, to replace REMARK.

Targets of Opportunity (ToO):
   Bernie McNamara from NMSU has had in the past used APO for ToO
programs.  There is definitely precedent for doing this kind of thing;
APO is well set up for it, although we don't have any formal policy in
place; when ToO programs arise, people simply contact all relevant
people (Gillespie, Turner, and the local APO schedulers) and make
arrangements as best they can.  At the moment, this is happening
infrequently enough that it is not causing any major perturbation on
the schedule.   We agreed that this practice should be encouraged, and
briefly discussed whether the observatory should be in the position of
actively carrying out the ToO programs (e.g., if someone writes a
proposal to take spectra of the next bright supernova to go off, then
the observatory would take responsibility for getting those spectra as
soon as they hear about a suitable supernova).  We realized that it
would be preferable to give the PI the responsibility to ask for the
time each time a relevant ToO comes up.  Of course, the way for these
PI's can be smoothed by writing a proposal ahead of time, as it were,
warning the Observatory and the Director that if a ToO comes up, they
will be asking for time.

   In the meantime, each User's Committee member will query the
members of their institution about interest in ToO programs.

  DIS grating:
  There has been interest in getting gratings for DIS of resolution
intermediate between the high and low resolution gratings currently in
place.  Alan Uomoto has been looking into this, and suggests a pair of
gratings at 300 and 600 l/mm, respectively.   To finalize this will
require knowing the details of DIS' optical prescription.  Alan plans
to buy something off the shelf; when he finds something appropriate,
he will present it to the User's Committee at the next user's
committee meeting.

  DIS chips:
   There is wide-spread agreement that the chips currently in the DIS
are about 2 generations too old.  They have a high read noise,
persistent ghosts, and a variety of other problems.  There is an
opportunity to get *free* CCDs through an NSF program through Lick
Observatory (2Kx2K, 2Kx4K Orbit and Loral chips), although these chips
come in wafer form, without any packaging.  In addition, they are
thick chips, and would need to be thinned.  Alan estimates ~$50,000
and 2 years development time would be required to do all that is
necessary to get them ready to do science. Full details of this
program, which requires writing a science proposal, can be found at: Proposals are due
soon (Feb 21)!
  Alan feels that the development cost and time to get these chips to
work is prohibitive, and is looking into the possibility of using
chips from the HST Advanced Camera (although those chips have not yet
appeared).  He asks whether, if financial push comes to shove, we
might delay the recoating of the DIS optics to fix the UV throughput
problem, in order to use the money ($5-7K saved) to put into the
chips.   This is not very much money, but it could make the
difference.  The User's Committee will query their users about this.

  Pointing problems:
   Since the shutdown, the telescope has been having a lot of pointing
glitches: things will work fine for a while, and then the pointing
will be *way* off, and/or images will be trailed.  20-30% of time on
those nights on which this is bad is lost.  It is believed that this
is due to falling behind on maintaining the lubrication on the azimuth
encoder..  It is a day of work, plus some nightime
effort, to clean these; this is currently planned for Feb 13 (the
first half of that night is engineering time).
   There are related lubrication and wear issues with the altitude
drives; this
does not appear to be a cause of the current pointing troubles.

  Recent Computer Glitches:
     There have been quite a few instances lately in which there has
been loss of communication between various of the systems which
operate the telescope and the instruments.  There are several
potential causes:
   1. We have upgrated to latest Solaris on MC; there may be some
subtle incompatibilities between machines.
   2. The system may be becoming sensitive to fragmented disks; all
the software could use a rebuilding (which is quite an extensive
process; perhaps best done when the observatory is shut down for
several days by heavy snow).
   3. Finally, it is possible that there are some problems with
various electrical connections.
   All these are being looked into.

  Monitoring the Telescope Throughput:
    As everyone knows, the realuminization of the primary, secondary,
and tertiary have dramatically increased the throughput of the
telescope.  See the report by Watson and Ledlow on  It is planned to repeat
the Watson-Ledlow throughput measurements with DIS periodically to
monitor this throughput.  This can be done in service mode, in
twilight (it remains a matter of some discussion how often this should
be done.  Weekly?  Monthly?  It of course requires photometric
   We obviously want to keep mirror surfaces pristine, which will be
done with use of cleaning with CO_2 snow, and a more agressive
avoidance of precipitation on the mirror.
  It would also be good to monitor throughput of SPICAM and the guider
as well, to isolate problems that might arise in these instruments.

  Instrument issues:
   DIS: the fiber optic recabling by Uomoto and Brinkmann is not yet
complete; they hope to finish this soon, which should eliminate a
source of AC pickup.

  GRIM: In good seeing, the focus is seen to be nonuniform over the
chip (apo35-grim #34); it is not clear whether this is a new
development, or has been true for a while.  It is possible that when
GRIM was first installed, it was adjusted to the out-of-collimation
telescope, and it is now tilted with respect to the focal plane now
that the telescope has been properly collimated.  This change of focus
is *not* seen with the other instrument, in particular the guider.

  CHAOS and the artificial guide star laser are being removed from the
telescope, for repairs.  The group needs more than the 2-3
nights/quarter that they are getting, and are planning to look for
another, presumably smaller, telescope, on which to do their
development work. 

  The polishing contract for the new secondary is close to getting bids.

  Last meeting's minutes are approved.

  Next meeting, March 10, 12:30 PM.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 126 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to